OpenStudy (anonymous):

A chef draws cookies randomly from a box containing 6 cookies of the same shape and size. There is 1 chocolate cookie, 3 almond cookies, and 2 butter cookies. He draws 1 cookie and then draws another cookie without replacing the first one. Find the probability of picking 1 almond cookie followed by another almond cookie, and show the equation used. ^@skielerlucas04 here's the first one.

OpenStudy (anonymous):

oops @Skielerlucas04 ^

OpenStudy (anonymous):

1 + 3 + 2 = 6 1 / 6 = 0.16 x 100 = 16% 1 / 5 = 0.2 x 122 = 20%

OpenStudy (anonymous):

heres the next question

OpenStudy (anonymous):

Minni is arranging 3 different music CDs in a row on a shelf. Create a sample space for the arrangement of a jazz CD (J), a pop CD (P), and a rock CD (R).

OpenStudy (anonymous):

3/6 * 2/5 = 6/30 =1/5

OpenStudy (anonymous):

thats the first problem^

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ya sorry

OpenStudy (anonymous):

no youre fine :)

OpenStudy (anonymous):

Minni is arranging 3 different music CDs in a row on a shelf. Create a sample space for the arrangement of a jazz CD (J), a pop CD (P), and a rock CD (R).

OpenStudy (anonymous):

that is question 2 @Skielerlucas04 ^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

J=4 P=3 R=5

OpenStudy (anonymous):

ok next question

OpenStudy (anonymous):

A box has 1 red marble, 3 blue marbles, and 4 green marbles. Maya draws a blue marble randomly from the box, replaces it, and then draws another marble randomly. What is the probability of drawing 2 blue marbles in a row? Explain your answer.

OpenStudy (anonymous):

"explain" -_-

OpenStudy (anonymous):

ik ik

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

this will take awhile

OpenStudy (anonymous):

ok take your time

OpenStudy (anonymous):

do you wanna do qa different problem and then go back to this

OpenStudy (anonymous):

at the first draw, the probability for getting a blue marble, is: 3/8 then Maya replaces it into the box at the second draw the probability to getting a blue marble, is the same, namely 3/8, so both probabilities are 3/8 then the requested probability is: 3/8 * 3/8 = 9/64

OpenStudy (anonymous):

nevermind lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Greg has 4 shirts: a white one, a black one, a red one, and a blue one. He also has two pairs of pants, one blue and one tan. What is the probability, if Greg gets dressed in the dark, that he winds up wearing the white shirt and tan pants? Show your work.

OpenStudy (anonymous):

that is the last one

OpenStudy (anonymous):

explain and show your work again :/

OpenStudy (anonymous):

Here the probability to have a white shirt is: 1/4 since there is one favorable case and four possible cases where as the probability to have a tan pants is: 1/2 since there is one favorable case and 2 possible cases so the probability of a white shirt is 1/4 and the probability of tan pants is ½. Then the requested probability is: 1/2 * 1/4 = 1/8

OpenStudy (anonymous):

ok thanks ill submit my test now

OpenStudy (anonymous):

YYYAAAAYYYY

OpenStudy (anonymous):

lol ill talk to you later i g2g bye thanks for the help ill kik you

OpenStudy (anonymous):

can we just chat on here?

OpenStudy (anonymous):

sure @Skielerlucas04